### v1

# Supplementary Materials

Abstract
In this study, we introduce a method for learning group (known or unknown) equivariant functions by learning the associated quadratic form $x^{T}Ax$ corresponding to the group from the data. Certain groups, known as orthogonal groups, preserve a specific quadratic form, and we leverage this property to uncover the underlying symmetry group under the assumption that it is orthogonal. By utilizing the corresponding unique symmetric matrix and its inherent diagonal form, we incorporate suitable inductive biases into the neural network architecture, leading to models that are both simplified and efficient. Our approach results in an invariant model that preserves norms, while the equivariant model is represented as a product of a norm-invariant model and a scale-invariant model, where the “product” refers to the group action.
Moreover, we extend our framework to a more general setting where the function acts on tuples of input vectors via a diagonal (or product) group action. In this extension, the equivariant function is decomposed into an angular component extracted solely from the normalized first vector and a scale-invariant component that depends on the full Gram matrix of the tuple. This decomposition captures the inter-dependencies between multiple inputs while preserving the underlying group symmetry.
We assess the effectiveness of our framework across multiple tasks, including polynomial regression, top quark tagging, and moment of inertia matrix prediction. Comparative analysis with baseline methods demonstrates that our model consistently excels in both discovering the underlying symmetry and efficiently learning the corresponding equivariant function.
Learning Equivariant Functions via Quadratic Forms
Pavan Karjol1          Vivek V Kashyap1          Rohan Kashyap2          Prathosh A P1
1Department of Electrical Communication Engineering, Indian Institute of Science, Bengaluru, Karnataka 2Computer Science, Carnegie Mellon University, Pittsburgh, Pennsylvania, USA
1 Introduction
Symmetry serves as a fundamental principle in modern deep learning, providing powerful inductive biases that enhance model generalization and sample efficiency Cohen (2013); Cohen and Welling (2014); Finzi et al. (2021); Dehmamy et al. (2021). When the symmetry group underlying a system is known a priori, this symmetry can be explicitly encoded into the neural network architecture through invariant or equivariant operations. This architectural integration ensures the model maintains equivariance with respect to the predefined group transformations present in the data manifold. The classical example is the Convolutional Neural Networks (CNNs), which maintain translational equivariance through weight sharing across spatial dimensions.
Recent works have explored two primary approaches to incorporating symmetries: augmenting training data using symmetry structures (Benton et al., 2020) and designing architectures with intrinsic symmetry properties (Zaheer et al., 2017; Kicki et al., 2020). (Benton et al., 2020) propose a data augmentation approach by parameterizing the data distribution while (Zhou et al., 2020) introduce a meta-learning framework through reparameterization of the network layers. In our work, we present a framework for learning group-equivariant functions through quadratic forms, emphasizing applications to orthogonal groups. Our analysis proves that these functions admit a canonical decomposition into two fundamental components: scale-invariant and norm-invariant terms.
Although substantial progress has been achieved in harnessing symmetries across various domains—including images (Anselmi et al., 2019; Kondor and Trivedi, 2018), graphs (Han et al., 2022), and dynamical systems such as partial differential equations (PDEs) (Horie and Mitsume, 2023)—significant challenges remain. A central difficulty in integrating these group symmetries within model architectures is the often-required explicit, a priori knowledge of the underlying symmetry. Furthermore, existing models frequently suffer from high computational demands, stemming from requirements such as null-space computations and implementation of exponential maps. This complexity limits their scalability and flexibility. Additionally, in many practical applications, the group action on the data may not be immediately apparent, making it difficult to effectively leverage these symmetries in current group-equivariant models.
Consequently, existing models are constrained in their ability to effectively manage complex group symmetries, especially within real-world data contexts. This limitation is similarly observed in dynamical systems, where simple symmetries in a low-dimensional state space can appear as highly non-linear transformations when projected into high-dimensional measurements. In this work, we propose a novel framework for learning equivariant functions, which integrates two invariant functions and is applicable regardless of whether the underlying symmetry group is known or unknown, given certain constraints.
Our approach focuses on orthogonal groups and their inherent geometric properties. Specifically, we are motivated by a fundamental property of the standard orthogonal group $O(n)$: its preservation of the Euclidean norm, which leads to orbits defined by distinct norm values. In addition, we also observe scale invariance in the case $O(n)$-equivariant functions. We demonstrate that this characteristic extends to other orthogonal groups preserving an arbitrary quadratic form, $x^{\top}Ax$, and we exploit these geometric correspondences to construct a systematic framework for symmetry discovery via neural networks. Furthermore, we explore the Lorentz group, which preserves the Minkowski norm and embodies the essential symmetry underlying the spacetime structure in physics. In localized regions of spacetime, where gravitational effects are negligible, physical laws exhibit Lorentz invariance in a manner consistent with special relativity.
In Section 4, we provide a comprehensive mathematical framework for our proposed method, particularly with respect to orthogonal groups, establishing its scalability and efficiency as a model for real-world applications.
1.1 Contributions
  • We establish that any $G$-equivariant function $f:X\rightarrow\mathbb{R}^{m}$, with $X\subset\mathbb{R}^{n}$ and $G$ being the orthogonal group preserving the quadratic form $x^{T}Ax$, can be represented as the product of two invariant functions—one that is norm-invariant and another that is scale-invariant—with respect to the underlying quadratic form. Here, the product is to be understood in the sense of the group action.
  • We extend this representation to functions defined on tuples of input vectors via a diagonal (product) group action on the product space $X_{p}=X\times\cdots\times X$. In this extended setting, the equivariant function is decomposed into an angular component—extracted solely from the normalized first vector—and a scale-invariant component that depends on the full Gram matrix of the tuple.
  • Leveraging this canonical decomposition into scale- and norm-invariant functions, we introduce a framework for symmetry discovery in cases where the underlying symmetry group is unknown. In this framework, uncovering the symmetry in the data corresponds to learning the appropriate quadratic form, which enables us to adaptively identify the symmetries present within a given dataset.
  • We validate our proposed method across several tasks that leverage invariance and equivariance properties. In particular, we demonstrate its effectiveness on polynomial regression, top quark tagging classification (which exploits Lorentz invariance), and the prediction of moment of inertia matrices (an $O(n)$-equivariant task).
2 Related Work
2.1 Group Equivariance
Significant progress has been made both theoretically and practically in symmetry learning for deep learning architectures (Kondor et al., 2018; Bekkers, 2021; Finzi et al., 2021; Bronstein et al., 2021; Dehmamy et al., 2021). The classical example is the G-equivariant neural networks (G-CNNs) (Cohen and Welling, 2014) which extends CNN’s to accommodate a broader class of symmetry groups beyond translations. Building on this, (Kondor and Trivedi, 2018) developed convolution formulations for arbitrary compact group actions, while (Cohen et al., 2019) further extended G-CNNs to homogeneous spaces through equivariant linear mappings.
Equivariant multi-layer perceptrons (EMLPs) for general groups were introduced by (Finzi et al., 2021), leveraging null space computations to achieve equivariance; however, scalability challenges persist, especially for higher-order groups and also requires explicit knowledge of the group to encode the group representations. Similarly, (Otto et al., 2023) proposed a linear-algebraic approach that enforces symmetries through linear constraints and convex penalties, though this approach also encounters scalability limitations.
In related work, (Park et al., 2022) employed symmetric embedding networks to learn equivariant features for downstream tasks, although their method requires prior knowledge of the symmetry group. Contrastive learning methods were also developed by (Shakerinava et al., 2022; Dangovski et al., 2022) for self-supervised equivariant representations. Symmetry learning has found practical applications in domains such as protein structure prediction, where molecular properties require invariance for 3D rotations (Yim et al., 2023).
2.2 Automatic Symmetry Discovery
Advances have also been made in the area of automatic symmetry discovery for neural networks. (Dehmamy et al., 2021) introduced Lie algebra convolution networks (L-Conv), achieving equivariance to continuous groups by leveraging Lie algebras. Complementary to this, (Benton et al., 2020) developed parameterized distributions over training data augmentations, while (Zhou et al., 2020) proposed a meta-learning approach that reparameterizes network layers to address symmetry adaptivity. Focusing on invariant representations, (Anselmi et al., 2019) explored regularization techniques for unsupervised orbit learning. (Yang et al., 2023) introduced LieGAN, a generative adversarial framework for symmetry discovery that leverages the principle of distribution preservation under group transformations.
More recently, (Yang et al., 2024) proposed Latent LieGAN, which approximates non-linear symmetries via non-linear group actions, particularly suited for high-dimensional dynamical systems. Additionally, (Karjol et al., 2023) introduced a structural decomposition approach for identifying locally symmetric, dihedral, and cyclic subgroups through multi-armed bandit algorithms.
3 Preliminaries
Definition 3.1 (Group).
A group $(G,\cdot)$ is a set G equipped with a binary operation $\cdot:G\times G\to G$ satisfying the following properties:
  • Closure: For all $a,b\in G$, $a\cdot b\in G$.
  • Associativity: For all $a,b,c\in G$, $(a\cdot b)\cdot c=a\cdot(b\cdot c)$.
  • Identity element: There exists an element $e\in G$ such that for every $a\in G$, $e\cdot a=a\cdot e=a$.
  • Inverse element: For each $a\in G$, there exists an element $b\in G$ (often denoted $a^{-1}$) such that $a\cdot b=b\cdot a=e$.
Definition 3.2 (Group Action).
A group action of a group $G$ on a set $X$ is a function $G\times X\to X,\quad(g,x)\mapsto g\cdot x,$ such that for all $g,h\in G$ and $x\in X$:
  • Identity: $e\cdot x=x$, where $e$ is the identity element of $G$.
  • Compatibility: $(g\cdot h)\cdot x=g\cdot(h\cdot x)$.
Definition 3.3 (Orbit).
Given a group action of $G$ on a set $X$ the orbit of an element $x\in X$ under $G$ is the set:
$G\cdot x=\{g\cdot x\mid g\in G\}.$ (1)
Definition 3.4 (General Linear Group).
The general linear group $\mathrm{GL}(n,\mathbb{R})$ is the group of all $n\times n$ invertible matrices with entries from real numbers with group operation given by matrix multiplication.
Definition 3.5 (Invariant Function).
An invariant function under a group action is a function $f:X\to Y$ that remains unchanged when the group acts on its input. Formally, if $G$ acts on $X$, then $f$ is invariant if $f(g\cdot x)=f(x)$ for all $g\in G$ and $x\in X$. Invariance often signifies that $f$ depends only on certain properties that are preserved by the group action.
Definition 3.6 (Equivariant Function).
An equivariant function between two sets with group actions is a function that respects the group action. If a group $G$ acts on both sets $X$ and $Y$, then a function $f:X\to Y$ is equivariant if $f(g\cdot x)=g\cdot f(x)$ for all $g\in G$ and $x\in X$. Equivariance means $f$ “commutes” with the group action and transforms in a manner consistent with the structure of $G$. Invariant function is a special case of Equivariant Function where $g\cdot f(x)=f(x),\ \quad\forall g\in G$.
Figure 1: Schematic representation of the proposed method. $\phi_{n}$ and $\phi_{s}$ denote neural networks, where the norm-invariant network takes input norms, and the scale-invariant network takes normalized inputs where $\|x\|_{A}=\operatorname{sign}(x^{T}Ax)\sqrt{|x^{T}Ax|}$.
4 Proposed Method
We present a general framework for learning group-equivariant functions by embedding appropriate quadratic forms within our network architecture. Specifically, we focus on groups that preserve quadratic forms of the type $x^{T}Ax$”, known as orthogonal groups. Our framework supports two cases: when the underlying group is known and when it is unknown. In the first case, we fix the quadratic form within the architecture. In the second case, we allow the quadratic form to be learned. We begin with a brief overview of quadratic forms before presenting our main contributions.
4.1 Quadratic form
Quadratic form is a type of mathematical expression in which a function of several variables involves terms of degree two (or order two). Let $A\in\mathbb{R}^{n\times n}$ be an $n\times n$ matrix and $x\in\mathbb{R}^{n}$. In general, a quadratic form in $x\in\mathbb{R}^{n}$ can be written as,
$q\mathopen{}\left({x;A}\,\negthickspace\right)\mathclose{}:=x^{T}Ax$ (2)
Each quadratic form defined on $\mathbb{R}^{n}$ is uniquely determined by a symmetric matrix. This is because, we can write,
$q(x;A)=x^{T}Ax=\frac{1}{2}\mathopen{}\left[{x^{T}\mathopen{}\left({A+A^{T}}\,\negthickspace\right)\mathclose{}x}\,\negthickspace\right]\mathclose{},\quad\forall x\in X$ (3)
Here, the $\frac{1}{2}\mathopen{}\left({A+A^{T}}\,\negthickspace\right)\mathclose{}$ is the unique symmetric matrix for the quadratic form $q\mathopen{}\left({.;A}\,\negthickspace\right)\mathclose{}$. Hence, we can simply assume that $A$ is symmetric matrix in any quadratic form $q(x;A)$.
4.1.1 Diagonalizing the Quadratic Form
Given a quadratic form $q(x;A)=x^{T}Ax$, where $A$ is an $n\times n$ symmetric matrix, we aim to simplify this form by transforming it into a diagonal form.
Since $A$ is symmetric, it can be diagonalized by an orthogonal matrix $U$. This means there exists an orthogonal matrix $U$ such that:
$A=U^{T}DU,$ (4)
where $D$ is a diagonal matrix containing the eigenvalues of $A$.
Now, we can express the quadratic form in terms of this diagonalization as follows:
$q\mathopen{}\left({x;A}\,\negthickspace\right)\mathclose{}=x^{T}\mathopen{}\left({U^{T}DU}\,\negthickspace\right)\mathclose{}x=\mathopen{}\left({Ux}\,\negthickspace\right)\mathclose{}^{T}D\mathopen{}\left({Ux}\,\negthickspace\right)\mathclose{}=y^{T}Dy,$ (5)
where $y=Ux$. Thus, learning the quadratic form $q(x;A)$ reduces to learning an orthogonal matrix $U$ and a diagonal matrix $D$.
4.2 Orthogonal Group: $O(n;q)$
Consider the set,
$G:=\mathopen{}\left\{{U\in GL\mathopen{}\left({n;\mathbb{R}}\,\negthickspace\right)\mathclose{}:U^{T}AU=A}\,\negthickspace\right\}\mathclose{},$ (6)
which can easily be shown to be a Lie subgroup of $GL(n;\mathbb{R})$. Such groups are called orthogonal groups, denoted as $O(n;q)$, and they preserve the quadratic form $q(x;A)=x^{T}Ax$, meaning that $q(x;A)=q(Rx;A)$ for all $R\in O(n;q)$.
Examples of orthogonal groups include the standard orthogonal group that preserves the Euclidean norm $x^{T}x$, and the Lorentz group, which preserves the Minkowski norm $x^{T}\eta x$ where $\eta=\operatorname{diag}\left(1,-1,-1,-1\right)$, among others.
In this work, we focus on learning a function $f:X\rightarrow\mathbb{R}^{m}$, where $f$ is equivariant under an orthogonal group $G$ that preserves a given or unknown quadratic form $q(x;A)$. In other words, $A$ may be known or unknown.
We now state our first result in the following theorem, which characterizes the set of orbits under the action of an orthogonal group $G$. Specifically, each $G$-orbit corresponds to a unique value of the quadratic form.
Theorem 4.1.
Let $A$ be a non-zero $n\times n$ matrix, and define $U:=\mathopen{}\left\{{x\in\mathbb{R}^{n}:x^{T}Ax=0}\,\negthickspace\right\}\mathclose{}$. Let $G$ be an orthogonal group that preserves the quadratic form $x^{T}Ax$, and consider the action of $G$ on $\mathbb{R}^{n}\setminus U$. For a given $c\in\mathbb{R}\setminus\{0\}$, define the set $\mathcal{O}_{c}:=\mathopen{}\left\{{x\in\mathbb{R}^{n}:x^{T}Ax=c}\,\negthickspace\right\}\mathclose{}$. Then, the set of orbits under the action of $G$ is given by
$\mathcal{O}\mathopen{}\left({G}\,\negthickspace\right)\mathclose{}=\mathopen{}\left\{{\mathcal{O}_{c}:c\in\mathbb{R}^{n}\setminus\{0\}\And\mathcal{O}_{c}\neq\varnothing}\,\negthickspace\right\}\mathclose{}.$ (7)
Using the above result, we show that any $G$-equivariant function can be expressed in a canonical form that combines a norm-invariant (equivalently, $G$-invariant) function and a scale-invariant function. This formulation is formalized in the following theorem.
Theorem 4.2.
Let $A$ be a non-zero $n\times n$ matrix, and define $U:=\mathopen{}\left\{{x\in\mathbb{R}^{n}:x^{T}Ax=0}\,\negthickspace\right\}\mathclose{}$. Any $G$-equivariant function $f:\mathbb{R}^{n}\setminus U\rightarrow\mathbb{R}^{m}$, where $G$ is an orthogonal group that preserves the quadratic form $x^{T}Ax$, can be expressed as,
$f(x)=\phi_{s}\mathopen{}\left({\frac{x}{\|x\|_{A}}}\,\negthickspace\right)\mathclose{}\cdot\phi_{n}\mathopen{}\left({\|x\|_{A}}\,\negthickspace\right)\mathclose{},$ (8)
for some functions $\phi_{s}:\mathbb{R}^{n}\rightarrow G$ and $\phi_{n}:\mathbb{R}\rightarrow\mathbb{R}^{m}$. Here, $\|z\|_{A}$ denotes the norm associated with the quadratic form $A$, defined by $\|z\|_{A}=\operatorname{sign}(z^{T}Az)\sqrt{|z^{T}Az|}$.
The scale invariance part of the function decomposition of (8) is illustrated in Figure 2.
Remark 4.1.
The scale invariance function $\phi_{s}$ in (8) is required to be $G$-equivariant. However, because its input is normalized with respect to the quadratic form defined by $A$, $\phi_{s}$ effectively operates on a single orbit, unlike the original function $f$ in (8) which acts on multiple orbits. Consequently, one could incorporate an additional inductive bias toward $G$-equivariance for $\phi_{s}$ (for example, by introducing a regularizer) beyond the inherent scale invariance. In our experiments, however, we observed that the built-in scale invariance suffices.
[…]
Figure 2: For an $O(2)$-equivariant function $f$, we have $f(x)=f\mathopen{}\left({R(\theta)r_{1}e_{1}}\,\negthickspace\right)\mathclose{}=\underbrace{R(\theta)}{}f\mathopen{}\left({r_{1}e_{1}}\,\negthickspace\right)\mathclose{}$. Similarly, $f(y)=\underbrace{R(\theta)}{}f\mathopen{}\left({r_{2}e_{1}}\,\negthickspace\right)\mathclose{};\>f(z)=\underbrace{R(\theta)}{}f\mathopen{}\left({r_{3}e_{1}}\,\negthickspace\right)\mathclose{}.$ Hence, $\phi_{s}(x)=\phi_{s}(y)=\phi_{s}(z)=R(\theta)$. This property holds for general orthogonal groups as well.
Remark 4.2.
In Theorem 4.2, the set $U$ has measure zero in $\mathbb{R}^{n}$. Therefore, including this set $U$ in our experiments has a negligible impact on the results, as it represents an insignificant portion of the space.
Next, we analyze the continuity and differentiability of the $G$-equivariant function as follows:
Proposition 4.1.
Consider Theorem 4.2. Suppose $A$ is either a positive or a negative definite matrix. Then, $\phi_{n}$ and $\phi_{s}$ are smooth ($C^{\infty}$) whenever $f$ is smooth ($C^{\infty}$). If $A$ is neither a positive nor a negative definite matrix, then $\phi_{n}$ and $\phi_{s}$ are continuous ($C^{0}$) whenever $f$ is continuous ($C^{0}$).
Please refer to the appendix for the proofs of all the Theorems and propositions.
4.3 Diagonal group action
We now generalize our approach to the setting where the function acts on tuples of vectors rather than on individual vectors. Specifically, let $X$ be a $G$-invariant subset of $\mathbb{R}^{n}\setminus U$, where
$U:=\{x\in\mathbb{R}^{n}:x^{T}Ax=0\},$
and define the product space
$X_{p}:=\underbrace{X\times X\times\cdots\times X}_{p\text{ times}}.$
On $X_{p}$, the group $G$ acts diagonally, that is,
$g\cdot(x_{1},x_{2},\ldots,x_{p})=(g\cdot x_{1},\,g\cdot x_{2},\,\ldots,\,g\cdot x_{p}).$
We recall the standard result (see Appendix, Classification of $p$-tuples up to the $A$-orthogonal group) which states that any two $p$-tuples of vectors in $\mathbb{R}^{n}$ that share the same Gram matrix (with respect to a non-degenerate symmetric bilinear form) are related by an isometry (Jacobson, 1985). Building upon this well-established fact, we derive a new decomposition theorem for functions equivariant under the diagonal $G$-action on tuples. This new result, formalized in the following theorem, extends our previous method to a broader and more general setting.
Theorem 4.3 (Extended Decomposition for $G$-Equivariant Functions on Tuples).
Let $A$ be a non-singular $n\times n$ real matrix and define Let
$G:=\{g\in\mathrm{GL}(n,\mathbb{R}):g^{T}A\,g=A\}$
be the orthogonal group preserving the quadratic form $x^{T}Ax$. Let $X\subset\mathbb{R}^{n}$ be a $G$-invariant set. For any positive integer $p$, consider the diagonal action of $G$ on
$X_{p}:=\underbrace{X\times X\times\cdots\times X}_{p\text{ times}},$
defined by
$g\cdot(x_{1},x_{2},\ldots,x_{p})=(g\cdot x_{1},\,g\cdot x_{2},\,\ldots,\,g\cdot x_{p}).$
Then every $G$-equivariant function
$f:X_{p}\to\mathbb{R}^{m}$
can be decomposed as
$f(x)=\phi_{s}\mathopen{}\left({\frac{x_{1}}{\|x_{1}\|_{A}}}\,\negthickspace\right)\mathclose{}\cdot\phi_{n}\mathopen{}\left({\operatorname{Gram}(x)}\,\negthickspace\right)\mathclose{},$ (9)
where $x=\mathopen{}\left({x_{1},\ldots,x_{p}}\,\negthickspace\right)\mathclose{}$ and Gram matrix
$\operatorname{Gram}(x_{1},\ldots,x_{p})=\begin{pmatrix}x_{1}^{T}Ax_{1}&\cdots&x_{1}^{T}Ax_{p}\\ \vdots&\ddots&\vdots\\ x_{p}^{T}Ax_{1}&\cdots&x_{p}^{T}Ax_{p}\end{pmatrix}$,
encodes all pairwise inner products with respect to the quadratic form defined by $A$.
Remark 4.3.
When AA is non-singular, Theorem 4.2 is a special case of Theorem 4.3. Notably, Theorem 4.2 remains applicable even when AA is singular. In contrast, Theorem 4.3 would require further modifications to accommodate the singular case, and we leave these extensions for future work.
4.4 Learning Equivariant Functions
Given data $\{\left(x_{i},y_{i}=f(x_{i})\right)\}_{i\in[m]}$, where $f$ is a $G$-equivariant function for some known or unknown orthogonal group $G$ that preserves the quadratic form $x^{T}A_{0}x$, we implement the functions $\phi_{s}$ and $\phi_{n}$ through neural networks. We denote these networks by the same notation, with $\phi_{s}$ and $\phi_{n}$ representing the scale-invariant and norm-invariant components of the function, respectively.
4.4.1 $G$ is Known
When the group $G$ is known, we fix the corresponding quadratic form in our network architecture, specifically fixing the matrix $A$. In this case, we aim to minimize the following loss function:
$\arg\min\limits_{\theta}\sum\limits_{i}\mathcal{L}\mathopen{}\left({y_{i},\phi_{s}\mathopen{}\left({\frac{x}{\|x\|_{A}}}\,\negthickspace\right)\mathclose{}\cdot\phi_{n}\mathopen{}\left({\|x\|_{A}}\,\negthickspace\right)\mathclose{}}\,\negthickspace\right)\mathclose{}$ (10)
where $\theta=\left\{\theta_{s},\theta_{n}\right\}$ are the weights of the networks corresponding to $\phi_{s}$ and $\phi_{n}$, respectively.
4.4.2 $G$ is Unknown: Symmetry Discovery Framework
In this case, the goal is to discover the underlying symmetry group $G$ by learning the appropriate matrix $A$. We seek to minimize the following loss function:
$\arg\min\limits_{A,\theta}\sum\limits_{i}\mathcal{L}\mathopen{}\left({y_{i},\phi_{s}\mathopen{}\left({\frac{x}{\|x\|_{A}}}\,\negthickspace\right)\mathclose{}\cdot\phi_{n}\mathopen{}\left({\|x\|_{A}}\,\negthickspace\right)\mathclose{}}\,\negthickspace\right)\mathclose{},$ (11)
Thus, discovering the underlying group $G$ corresponds to learning the appropriate symmetric matrix $A$. However, as discussed earlier, this problem reduces to learning the appropriate orthonormal matrix $U$ and diagonal matrix $D$ due to the diagonalization of the symmetric matrix $A$. Therefore, we can rewrite the objective function in Eq. (11) as:
$\arg\min\limits_{U,D,\theta}\sum\limits_{i}\mathcal{L}\mathopen{}\left({y_{i},\phi_{s}\mathopen{}\left({\frac{\hat{x}}{\|\hat{x}\|_{D}}}\,\negthickspace\right)\mathclose{}\cdot\phi_{n}\mathopen{}\left({\|\hat{x}\|_{D}}\,\negthickspace\right)\mathclose{}}\,\negthickspace\right)\mathclose{},$ (12)
where $\hat{x}=Ux$, $U\in O(n)$ and $D$ is a diagonal matrix. In this formulation, $U$ captures the orthogonal transformation and $D$ scales the components of the input $x_{i}$, thereby enabling the discovery of the underlying symmetry group.
For invariance tasks, we use only the $\phi_{n}$ network and omit $\phi_{s}$. When the group acts on the input domain via a product action, we adopt the function decomposition from (9). This decomposition is applied both when the group is explicitly known—where we incorporate it into (10)—and when the group is unknown, in which case it is used in (11).
Method Orthogonal Lorentz Arbitrary
$G$-Ortho-Nets $1.12\times 10^{-6}$ $8.72\times 10^{-7}$ $6.64\times 10^{-7}$
EMLP $6.64\times 10^{-2}$ $4.66\times 10^{-2}$ N/A
Simple MLP $7.83\times 10^{-5}$ $1.83\times 10^{-5}$ $2.41\times 10^{-5}$
Table 1: MSE on validation dataset using G-Ortho Nets is evaluated for Orthogonal $O(n)$, Lorentz $O(1,3)$, and Arbitrary Group for invariance task. To ensure a rigorous comparative analysis, all models are designed with an identical parameter count and trained for the same number of epochs, ensuring a fair and consistent evaluation. Note: EMLP failed to converge for the Arbitrary Group.
Method Moment of Inertia Top Tagging
Val Loss Accuracy AUROC
$G$-Ortho-Nets $1.16\times 10^{-5}$ $0.915$ $0.969$
EMLP $1.36\times 10^{-3}$ N/A N/A
LieGNN N/A $0.938$ $0.984$
Table 2: Performance Comparison on Moment of Inertia Prediction and Top Tagging Classification Tasks: The table shows that while G-Ortho Nets achieves slightly lower accuracy than LieGNN, it unifies symmetry discovery and classification in a single network. In contrast, LieGNN employs a two-stage approach using LieGAN for symmetry discovery followed by a separate GNN for classification.
Task $G$-Ortho-Nets LieGAN
$G$-Invariant Polynomial Regression $9.99$ $9.99$
Top Tagging $9.99$ $9.99$
Equivariant Polynomial Regression $9.99$ $9.99$
Moment of Inertia $9.99$ $9.99$
Table 3: Comparison of cosine similarity between $G$-Ortho-Nets and LieGAN across multiple tasks.
(a) GTGT
(b) LieGANLieGAN
(c) Ours
(d) GTGT
(e) LieGANLieGAN
(f) Ours
Figure 3: Comparison of ground truth and learned matrices. (Top) Top quark tagging: the ground truth matrix $A=\operatorname{diag}(1,-1,-1,-1)$ is compared with matrices learned by LieGAN and $G$-Ortho-Nets . (Bottom) Invariant polynomial regression with an arbitrary matrix AA: the ground truth is compared with matrices recovered by LieGAN and $G$-Ortho-Nets. Both methods accurately recover the underlying matrix structure.
5 Experiments and Results
We evaluate our proposed framework on both $G$-invariant and $G$-equivariant tasks, where $G$ is an orthogonal group as defined in Eq. (6). To assess its performance, we consider three distinct types of groups: the standard orthogonal group, the Lorentz group, and a group that preserves a random quadratic form $x^{T}Ax$, where $A$ is a randomly generated symmetric matrix. Across all experiments, we compare the following model architectures:
  • G-Ortho-Nets: Our proposed method, which incorporates the product of a norm-invariant and a scale-invariant function.
  • EMLP: EMLP architecture with G-equivariant linear and bilinear layers as discussed in (Finzi et al., 2021).
  • LieGAN: Generative Adversarial Symmetry Discovery proposed by (Yang et al., 2023)
  • LieGNN: LieGAN integrated with a graph neural network.
  • Simple MLP: A standard feed-forward multilayer perceptron.
The experimental setup is detailed in the subsequent subsections.
5.1 $G$-Invariant Polynomial Regression
(a) GTGT
(b) LieGANLieGAN
(c) Ours
Figure 4: Comparison of the ground truth matrix AA (first plot) with matrices learned by LieGAN (second plot) and $G$-Ortho-Nets (third plot) for the $G$-Equivariant Polynomial Regression task.
We conduct a series of experiments to evaluate our method’s ability to learn invariant polynomial functions of the form:
$f(x)=\sin^{2}(x^{T}Ax)+\sin(x^{T}Ax)$ (13)
where the invariance structure is determined by the matrix $A$. We examine three distinct settings, each characterized by a different choice of $A$, to demonstrate the versatility and robustness of our approach.
Experimental Setting. We evaluate our method on three group-invariant polynomial regression tasks. For the orthogonal group, we set $A=I_{4}$ to construct an $O(4)$-invariant task. For the Lorentz group, we employ $A=\text{diag}(1,-1,-1,-1)$, corresponding to the Minkowski metric. For the arbitrary group case, we use a randomly chosen matrix $A$ to demonstrate the methods ability to generalize beyond well-studied groups in the literature. All tasks use the same target polynomial function.
Results. Table LABEL:tab:1 shows that $G$-Ortho Nets achieves lower MSE on the validation dataset compared to EMLP and Simple MLP baselines. Figures LABEL:fig:orthogonal_matrix6, and LABEL:fig:arbitrary_matrix illustrate that our method successfully recovers the underlying invariant structure, with learned matrices closely matching the ground truth across all three group settings. These results demonstrate the effectiveness of our method in both structure discovery and function approximation tasks.
5.2 Lorentz Invariance: Top Tagging
Experimental Setting. We evaluate our method on Lorentz-invariant top tagging task using the Top Quark Jet Dataset (Kasieczka et al., 2019). The task requires classifying jets originating from top quarks versus those from lighter quarks while preserving Lorentz symmetry. The underlying Lorentz metric $A=\text{diag}(1,-1,-1,-1)$ governs the symmetry structure, which our method discovers jointly with the classification task.
Results. $G$-Ortho Nets achieves comparable performance to LieGAN on the top tagging task, as shown in Table 2, with similar accuracy and AUROC scores. Additionally, Figure 6 demonstrates that our method successfully recovers the underlying Lorentz metric structure $A=\text{diag}(1,-1,-1,-1)$.
5.3 Equivariant Polynomial Regression
Experimental Setting. We evaluate our method on $G$-equivariant polynomial regression task with arbitrary group structure defined by matrix $A$. The target polynomial is
$f(x)=9(xx^{T}A)^{2}+2(xx^{T}A),$ (14)
which is equivariant with respect to $G$, since $f\mathopen{}\left({Ux}\,\negthickspace\right)\mathclose{}=Uf(x)U^{-1},\;\forall U\in G$.
Results. Figure LABEL:fig:equivariant_matrix compares the ground truth matrix $A$ with reconstructions from LieGAN and $G$-Ortho Nets. Table LABEL:tab:1 shows that both methods accurately recover the underlying group structure based on cosine similarity. However, $G$-Ortho Nets integrates symmetry discovery with equivariant function learning in a single framework, whereas LieGAN performs only structure discovery. This enables our method to simultaneously recover the group structure and learn the target polynomial function.
5.4 $O(3)$-Equivariance: Moment of Inertia
Experimental Setting. We evaluate our method on predicting the moment of inertia $\mathcal{I}$ from particle masses and positions $p=\{(x_{i},m_{i})\}_{i\in[M]}$. The moment of inertia is computed as
$\mathcal{I}=\sum_{i}m_{i}\left(x_{i}^{T}x_{i}\,I-x_{i}x_{i}^{T}\right),$ (15)
which is $O(3)$-equivariant. We compare our method against EMLP (Finzi et al., 2021), which requires known group structure, and the symmetry discovery method from (Yang et al., 2023).
Results. Figure LABEL:fig:moi_matrix compares the ground truth matrix $A=I_{3}$ (normalized) with reconstructions from LieGAN and $G$-Ortho-Nets. Both methods successfully recover the $O(3)$-equivariant structure. Table 2 shows that $G$-Ortho-Nets achieves lower MSE than EMLP on the prediction task. Table LABEL:tab:1 reports cosine similarity of 0.99 for both LieGAN and $G$-Ortho-Nets, indicating accurate structure recovery.
5.5 Limitations
Our method is currently restricted to generalized orthogonal groups and supports only standard group actions and diagonal (product) group actions. In future work, we aim to overcome these limitations by extending our framework to accommodate a broader range of symmetry groups and group actions.
6 Conclusion
In this work, we introduced a framework for learning group-equivariant functions using quadratic forms, with a focus on orthogonal groups. We showed that such functions can be expressed in a canonical form, incorporating both scale-invariant and norm-invariant components. Additionally, we extended our approach to a more general setting where the function acts on tuples of input vectors under a diagonal (or product) group action. In this extension, the equivariant function is decomposed into an angular component derived from the normalized first vector and a scale-invariant component based on the full Gram matrix of the tuple. This decomposition effectively captures inter-dependencies between multiple inputs while maintaining group symmetry.
Through extensive experiments, we demonstrated the ability of our method to discover underlying symmetries and efficiently learn equivariant functions across various tasks. In future work, we aim to extend our approach to broader symmetry groups beyond the general orthogonal group and explore a wider range of group actions.
Appendix A Appendix
(a) GTGT
(b) LieGANLieGAN
(c) $G$-Ortho-Nets
(d) GTGT
(e) LieGANLieGAN
(f) $G$-Ortho-Nets
Figure 5: Comparison of ground truth and learned matrices. Top $O(4)$-Invariant Polynomial Regression: The ground truth matrix $A=I_{4}$ is compared with matrices learned by LieGAN and $G$-Ortho-Nets. Both methods accurately recover AA (matching up to the first decimal) while modeling the polynomial $\sin^{2}(x^{T}Ax)+\sin(x^{T}Ax)$. Bottom $O(3)$-Equivariant Moment of Inertia: The ground truth matrix $A=I_{3}$ is compared with matrices learned by LieGAN and $G$-Ortho-Nets.
(a) GTGT
(b) LieGANLieGAN
(c) $G$-Ortho-Nets
Figure 6: Lorentz-Invariant Polynomial Regression: Ground truth $A=\operatorname{diag}(1,-1,-1,-1)$ (first plot) is compared with matrices learned by LieGAN (second plot) and $G$-Ortho-Nets (third plot). Both methods approximately recover AA and model the polynomial $\sin^{2}(x^{T}Ax)+\sin(x^{T}Ax)$.
You may include other additional sections here.
Appendix B Additional Results and Complete Proofs
Additional theoretical results and the complete proof of Theorem. 4.1 are provided here including supporting Lemmas.
Lemma B.1.
Let $D\in\mathbb{R}^{n\times n}$ be a diagonal matrix defined as:
$D=\sum_{i=1}^{n}d_{i}E_{i,i},$ (16)
where $d_{1},\dots,d_{n}>0$, and $E_{i,j}=e_{i}e_{j}^{T}$ with $e_{i}$ denoting the standard basis vectors in $\mathbb{R}^{n}$. Let $x=\sum_{i=1}^{n}x_{i}e_{i}\in\mathbb{R}^{n}$. Let $G$ be the orthogonal group preserving the quadratic form $x^{T}Dx$. Then, there exists $A\in G$ such that:
$Ax=\alpha e_{1},$ (17)
where $\alpha$ is an appropriate scaling factor such that $x^{T}Dx=x^{T}A^{T}DAx$.
Proof.
Define $S=\sum_{i=1}^{n}\sqrt{d_{i}}e_{i}e_{i}^{T}$. Then, we have:
$S^{-1}DS^{-1}=I,$ (18)
where $I$ is the identity matrix.
Consider the standard orthogonal group:
$O(n)=\mathopen{}\left\{{R\in GL\mathopen{}\left({n,\mathbb{R}}\,\negthickspace\right)\mathclose{}:R^{T}R=I}\,\negthickspace\right\}\mathclose{}.$
We know that if $y^{T}y=z^{T}z$ for any $y,z\in\mathbb{R}^{n}$, then there exists $R\in O(n)$ such that:
$Ry=z.$
Applying this result, we find $R\in O(n)$ such that:
$RSx=\beta e_{1},$ (19)
where $\beta$ is a scaling factor ensuring that the Euclidean norms of $Sx$ and $\beta e_{1}$ are equal.
Now, define $A=S^{-1}RS$. Since $S^{-1}$ is also a diagonal matrix, we compute:
$Ax$ $=S^{-1}RSx$
$=S^{-1}\beta e_{1}$
$=\alpha e_{1},$ (20)
where $\alpha=\frac{\beta}{\sqrt{d_{1}}}$ is an appropriate scaling factor.
Next, we show that $A\in G$. Consider $A^{T}DA$:
$A^{T}DA$ $=\mathopen{}\left({S^{-1}RS}\,\negthickspace\right)\mathclose{}^{T}D\mathopen{}\left({S^{-1}RS}\,\negthickspace\right)\mathclose{}$
$=S^{T}R^{T}\mathopen{}\left({S^{-1}DS^{-1}}\,\negthickspace\right)\mathclose{}RS$
$=S^{T}R^{T}IRS$
$=S^{T}IS$
$=D.$ (21)
Thus, $A\in G$, as required. ∎
Remark B.1.
The Lemma. B.1 is also applicable when $d_{i}<0,\;\forall i\in[n]$. We just need to appropriately modify the matrix $S$.
Lemma B.2.
Let $D\in\mathbb{R}^{n\times n}$ be a diagonal matrix defined as:
$D=\sum\limits_{i=1}^{k-1}d_{i}E_{i,i}\;\;:$ (22)
where $d_{1},\dots,d_{k-1}>0$, and $E_{i,j}=e_{i}e_{j}^{T}$ with $e_{i}$ denoting the standard basis vectors in $\mathbb{R}^{n}$. Consider a vector $y\in\mathbb{R}^{n}$ of the form:
$y=ae_{1}+be_{k},$
where $a,b\in\mathbb{R}$. Let $q(x;D)=x^{T}Dx$ be the quadratic form associated with $D$, and let $G=O(n,q)$ denote the orthogonal group that preserves this quadratic form, i.e.,
$G=\{R\in GL(n,\mathbb{R}):R^{T}DR=D\}.$
Then, there exists a matrix $A\in G$ such that:
$Ay=ae_{1}.$ (23)
Proof.
We construct $A$ explicitly and verify the required properties. Define $R\in\mathbb{R}^{n\times n}$ as:
$R=\sqrt{d_{1}}ab\,E_{k,1}-d_{1}a^{2}\,E_{k,k}+\sum_{i\in[n],\;i\neq 1,\;i\neq{k}}E_{i,i}.$ (24)
Define $S\in\mathbb{R}^{n\times n}$ as:
$S=\sqrt{d_{1}}E_{1,1}+\sum_{i\in[n],\;i\neq 1}E_{i,i}.$ (25)
Then:
$S^{-1}=\frac{1}{\sqrt{d_{1}}}E_{1,1}+\sum_{i\in[n],\;i\neq 1}E_{i,i}.$
Let $A=S^{-1}RS$. We claim that $A^{T}DA=D$ and $Ay=ae_{1}$
Step 1: Show $A^{T}DA=D$.
Using $A=S^{-1}RS$, we compute:
$A^{T}DA$ $=SR^{T}S^{-1}DS^{-1}RS$
First, compute $S^{-1}DS^{-1}$:
$S^{-1}DS^{-1}$ $=\mathopen{}\left({\frac{1}{\sqrt{d_{1}}}E_{1,1}+\sum_{i\in[n],\;i\neq 1}E_{i,i}}\,\negthickspace\right)\mathclose{}\mathopen{}\left({\sum\limits_{i=1}^{k-1}d_{i}E_{i,i}}\,\negthickspace\right)\mathclose{}\mathopen{}\left({\frac{1}{\sqrt{d_{1}}}E_{1,1}+\sum_{i\in[n],\;i\neq 1}E_{i,i}}\,\negthickspace\right)\mathclose{}$
$=E_{1,1}+\sum\limits_{i=2}^{k-1}d_{i}E_{i,i}.$ (26)
Consider $RS=\left(SR^{T}\right)^{T}$,
$RS$ $=\mathopen{}\left({\sqrt{d_{1}}ab\,E_{k,1}-d_{1}a^{2}\,E_{k,k}+\sum_{i\in[n],\;i\neq k}E_{i,i}}\,\negthickspace\right)\mathclose{}\mathopen{}\left({\sqrt{d_{1}}E_{1,1}+\sum_{i\in[n],\;i\neq 1}E_{i,i}}\,\negthickspace\right)\mathclose{}$
$=d_{1}abE_{k,1}+\sqrt{d_{1}}E_{1,1}-d_{1}a^{2}E_{k,k}+\sum_{i\in[n],\;i\neq 1,\;i\neq k}E_{i,i}.$ (27)
Thus, we have,
$A^{T}DA$ $=\mathopen{}\left({SR^{T}}\,\negthickspace\right)\mathclose{}\mathopen{}\left({S^{-1}DS^{-1}}\,\negthickspace\right)\mathclose{}\mathopen{}\left({RS}\,\negthickspace\right)\mathclose{}$
$=\mathopen{}\left({d_{1}abE_{1,k}+\sqrt{d_{1}}E_{1,1}-d_{1}a^{2}E_{k,k}+\sum_{i\in[n],\;i\neq 1,\;i\neq k}E_{i,i}}\,\negthickspace\right)\mathclose{}\mathopen{}\left({E_{1,1}+\sum\limits_{i=2}^{k-1}d_{i}E_{i,i}}\,\negthickspace\right)\mathclose{}\mathopen{}\left({RS}\,\negthickspace\right)\mathclose{}$
$=\mathopen{}\left({\sqrt{d_{1}}E_{1,1}+\sum\limits_{i=2}^{k-1}d_{i}E_{i,i}}\,\negthickspace\right)\mathclose{}\mathopen{}\left({RS}\,\negthickspace\right)\mathclose{}$
$=\mathopen{}\left({\sqrt{d_{1}}E_{1,1}+\sum\limits_{i=2}^{k-1}d_{i}E_{i,i}}\,\negthickspace\right)\mathclose{}\mathopen{}\left({d_{1}abE_{k,1}+\sqrt{d_{1}}E_{1,1}-d_{1}a^{2}E_{k,k}+\sum_{i\in[n],\;i\neq 1,\;i\neq k}E_{i,i}}\,\negthickspace\right)\mathclose{}$
$=d_{1}E_{1,1}+\sum\limits_{i=2}^{k-1}d_{i}E_{i,i}$
$=\sum\limits_{i=1}^{k-1}d_{i}E_{i,i}$
$=D$ (28)
Consider $A=S^{-1}RS$
$A$ $=S^{-1}\mathopen{}\left({RS}\,\negthickspace\right)\mathclose{}$
$=S^{-1}\mathopen{}\left({d_{1}abE_{k,1}+\sqrt{d_{1}}E_{1,1}-d_{1}a^{2}E_{k,k}+\sum_{i\in[n],\;i\neq 1,\;i\neq k}E_{i,i}}\,\negthickspace\right)\mathclose{}$
$=\mathopen{}\left({\frac{1}{\sqrt{d_{1}}}E_{1,1}+\sum_{i\in[n],\;i\neq 1}E_{i,i}}\,\negthickspace\right)\mathclose{}\mathopen{}\left({d_{1}abE_{k,1}+\sqrt{d_{1}}E_{1,1}-d_{1}a^{2}E_{k,k}+\sum_{i\in[n],\;i\neq 1,\;i\neq k}E_{i,i}}\,\negthickspace\right)\mathclose{}$
$=E_{1,1}+d_{1}abE_{k,1}-d_{1}a^{2}E_{k,k}+\sum_{i\in[n],\;i\neq 1,\;i\neq k}E_{i,i}.$ (29)
Step 2: Verify $Ay=ae_{1}$.
Substituting $y=ae_{1}+be_{k}$ into $Ay$, we have:
$Ay$ $=\mathopen{}\left({E_{1,1}+d_{1}abE_{k,1}-d_{1}a^{2}E_{k,k}+\sum_{i\in[n],\;i\neq 1,\;i\neq k}E_{i,i}}\,\negthickspace\right)\mathclose{}\mathopen{}\left({ae_{1}+be_{k}}\,\negthickspace\right)\mathclose{}$
$=ae_{1}+d_{1}a^{2}be_{k}-d_{1}a^{2}be_{k}$
$=ae_{1}.$ (30)
Thus, $Ay=ae_{1}$, as required. Conclusion: We have constructed $A\in G$ such that $Ay=ae_{1}$. Therefore, the lemma is proved. ∎
Remark B.2.
The Lemma. B.2 is also applicable when $d_{i}<0,\;\forall i\in\{1,\dots,k\}$. We just need to appropriately modify the matrices $S$ and $R$.
Lemma B.3.
Let $D\in\mathbb{R}^{n\times n}$ be a diagonal matrix defined as:
$D=\sum_{i=1}^{k-1}d_{i}E_{i,i}-\sum_{i=k}^{l}d_{i}E_{i,i},$ (31)
where $d_{1},\dots,d_{l}>0$, $k<l$, and $e_{i}$ denotes the standard basis vectors in $\mathbb{R}^{n}$. Let $x=ae_{1}+be_{k}\in\mathbb{R}^{n}$. Let $G$ be the orthogonal group preserving the quadratic form $x^{T}Dx$. Then, there exists $A\in G$ such that:
$Ax=\begin{cases}\alpha e_{1},&x^{T}Dx>0,\\ \alpha e_{k},&x^{T}Dx<0,\end{cases}$ (32)
where $\alpha$ is an appropriate scaling factor such that $x^{T}Dx=x^{T}A^{T}DAx$.
Proof.
We will prove the result for the case $x^{T}Dx>0$. Similar steps follow for the other case. The value of $x^{T}Dx$ is:
$x^{T}Dx=d_{1}a^{2}-d_{k}b^{2}.$ (33)
Since $d_{1}a^{2}-d_{k}b^{2}>0$, define the scalar $\beta$ as:
$\beta=\frac{1}{\sqrt{d_{1}a^{2}-d_{k}b^{2}}}.$ (34)
Define an $n\times n$ matrix $S$ as:
$S=\sqrt{d_{1}}E_{1,1}+\sqrt{d_{k}}E_{k,k}+\sum_{i\in[n],\;i\neq 1,k}E_{i,i}.$ (35)
Now define another $n\times n$ matrix $R$ as:
$R=\beta\mathopen{}\left({\sqrt{d_{1}}aE_{1,1}-\sqrt{d_{k}}bE_{1,k}-\sqrt{d_{k}}bE_{k,1}+\sqrt{d_{1}}aE_{k,k}+\sum_{i\in[n],\;i\neq 1,k}E_{i,i}}\,\negthickspace\right)\mathclose{}.$ (36)
Set $A=S^{-1}RS$. Our goal is to show that $A^{T}DA=D$ and $Ax=\alpha e_{1}$.
Step 1: Show that $A^{T}DA=D$.
Consider $A^{T}DA$:
$A^{T}DA$ $=\mathopen{}\left({SR^{T}S^{-1}}\,\negthickspace\right)\mathclose{}D\mathopen{}\left({S^{-1}RS}\,\negthickspace\right)\mathclose{}$
$=\mathopen{}\left({RS}\,\negthickspace\right)\mathclose{}^{T}\mathopen{}\left({S^{-1}DS^{-1}}\,\negthickspace\right)\mathclose{}\mathopen{}\left({RS}\,\negthickspace\right)\mathclose{}.$ (37)
We note that:
$S^{-1}DS^{-1}=E_{1,1}-E_{k,k}+\sum_{i=2}^{k-1}d_{i}E_{i,i}-\sum_{i=k+1}^{l}d_{i}E_{i,i},$ (38)
and:
$RS=\beta\mathopen{}\left({d_{1}aE_{1,1}-d_{k}bE_{1,k}-\sqrt{d_{1}d_{k}}bE_{k,1}+\sqrt{d_{1}d_{k}}aE_{k,k}+\sum_{i\in[n],\;i\neq 1,k}E_{i,i}}\,\negthickspace\right)\mathclose{}.$ (39)
Substituting (38) and (39) into (37), we get:
$A^{T}DA$ $=\beta^{2}\mathopen{}\left[{\frac{1}{\beta^{2}}d_{1}E_{1,1}-\frac{1}{\beta^{2}}d_{k}E_{k,k}+\sum_{i=2}^{k-1}d_{i}E_{i,i}-\sum_{i=k+1}^{l}d_{i}E_{i,i}}\,\negthickspace\right]\mathclose{}$
$=\sum_{i=1}^{k-1}d_{i}E_{i,i}-\sum_{i=k}^{l}d_{i}E_{i,i}$
$=D.$ (40)
Step 2: Show that $Ax=\alpha e_{1}$.
From (39) and (35), we compute:
$A$ $=S^{-1}RS$
$=\beta\mathopen{}\left({\sqrt{d_{1}}aE_{1,1}-\frac{d_{k}}{\sqrt{d_{1}}}bE_{1,k}-\sqrt{d_{1}}bE_{k,1}+\sqrt{d_{1}}aE_{k,k}+\sum_{i\in[n],\;i\neq 1,k}E_{i,i}}\,\negthickspace\right)\mathclose{}.$ (41)
Thus, the value of $Ax$ is:
$Ax$ $=\beta\mathopen{}\left[{\mathopen{}\left({\sqrt{d_{1}}a^{2}-\frac{d_{k}}{\sqrt{d_{1}}}b^{2}}\,\negthickspace\right)\mathclose{}e_{1}}\,\negthickspace\right]\mathclose{}$
$=\frac{\beta}{\sqrt{d_{1}}}\mathopen{}\left[{\mathopen{}\left({d_{1}a^{2}-d_{k}b^{2}}\,\negthickspace\right)\mathclose{}e_{1}}\,\negthickspace\right]\mathclose{}$
$=\frac{\beta}{\sqrt{d_{1}}}\mathopen{}\left[{\frac{1}{\beta^{2}}e_{1}}\,\negthickspace\right]\mathclose{}$
$=\frac{1}{\beta\sqrt{d_{1}}}e_{1}.$ (42)
Hence, $Ax=\alpha e_{1}$, where $\alpha=\frac{1}{\beta\sqrt{d_{1}}}=\frac{\sqrt{d_{1}a^{2}-d_{k}b^{2}}}{\sqrt{d_{1}}}$. ∎
Using the Lemma. B.1, B.2 and B.3, we provide the complete proof of Theorem. 4.1. See 4.1
Proof.
We have to show that each set $\mathcal{O}_{c}$ is a unique orbit. Suppose $x\in\mathcal{O}_{c}$, then by definition of orbit under the action of given orthogonal group, the orbif of $x$ is subset of $\mathcal{O}_{c}$, i.e., $G\cdot x\subset\mathcal{O}_{c}$. Now, we need to show the subset relation in reverse direction, i.e., $\mathcal{O}_{c}\subset G\cdot x$. This is equivalent to proving : if $x^{T}Ax=y^{T}Ay$, then $x$ and $y$ must lie in the same orbit.
To achieve this, we first diagonalize the symmetric matrix $A$. Let $UAU^{T}=D$, where $D$ is a diagonal matrix. By selecting an appropriate permutation matrix $P$, we further transform $D$ into a canonical form which partitions $D$ into the positive, negative, and zero eigenvalues of $A$. Without loss of generality, we assume this structure for $A$ and proceed with the proof.
Let $x\in\mathbb{R}^{n}$. We partition the diagonal matrix $D$ into positive, negative, and zero components:
$D=\mathopen{}\left[{\begin{array}[]{ccc}D_{p}&&\\ &D_{n}&\\ &&D_{z}\end{array}}\,\negthickspace\right]\mathclose{},$
where $D_{z}=\boldsymbol{0}$ (zero matrix). Correspondingly, partition $x$ as:
$x=\mathopen{}\left[{\begin{array}[]{c}x_{p}\\ x_{n}\\ x_{z}\end{array}}\,\negthickspace\right]\mathclose{},$
so that
$\|x\|_{D}^{2}$ $=\mathopen{}\left({x^{T}Dx}\,\negthickspace\right)\mathclose{}$
$=\mathopen{}\left({x_{p}^{T}D_{p}x_{p}}\,\negthickspace\right)\mathclose{}+\mathopen{}\left({x_{n}^{T}D_{n}x_{n}}\,\negthickspace\right)\mathclose{}+\mathopen{}\left({x_{z}^{T}D_{z}x_{z}}\,\negthickspace\right)\mathclose{}$
$=\|x_{p}\|_{D_{p}}^{2}+\|x_{n}\|_{D_{n}}^{2}+\|x_{z}\|_{D_{z}}^{2}.$ (43)
Since $\|x_{z}\|_{D_{z}}^{2}=0$ because $D_{z}=\boldsymbol{0}$, we have $\mathopen{}\left({x_{z}^{T}D_{z}x_{z}}\,\negthickspace\right)\mathclose{}=0$.
Step 1: Positive and Negative diagonal entries
Apply Lemma B.1 to $\mathopen{}\left({D_{p},x_{p}}\,\negthickspace\right)\mathclose{}$ and $\mathopen{}\left({D_{n},x_{n}}\,\negthickspace\right)\mathclose{}$, yielding matrices $A_{p}$ and $A_{n}$ such that:
$A_{p}^{T}D_{p}A_{p}=D_{p},\quad A_{p}x_{p}=\alpha e_{1},$
and
$A_{n}^{T}D_{n}A_{n}=D_{n},\quad A_{n}x_{n}=\beta_{1}e_{1}.$
Step 2: Zero diagonal entries
Since $\|x_{z}\|_{D_{z}}^{2}=0$, for any $y\in\mathbb{R}^{n-k-l}$, we can find an orthonormal matrix $A_{z}$ (i.e., $A_{z}^{T}A_{z}=I$) such that:
$A_{z}^{T}D_{z}A_{z}=D_{z}=\boldsymbol{0},\quad A_{z}x_{z}=\beta_{2}e_{1}.$
Define $A_{1}$ as:
$A_{1}=\mathopen{}\left[{\begin{array}[]{ccc}A_{p}&&\\ &A_{n}&\\ &&A_{z}\end{array}}\,\negthickspace\right]\mathclose{}.$
Since $A_{1}^{T}DA_{1}=D$, we have $A_{1}\in G$. Moreover,
$A_{1}x=\mathopen{}\left[{\begin{array}[]{ccc}A_{p}&&\\ &A_{n}&\\ &&A_{z}\end{array}}\,\negthickspace\right]\mathclose{}\mathopen{}\left[{\begin{array}[]{c}x_{p}\\ x_{n}\\ x_{z}\end{array}}\,\negthickspace\right]\mathclose{}=\mathopen{}\left[{\begin{array}[]{c}A_{p}x_{p}\\ A_{n}x_{n}\\ A_{z}x_{z}\end{array}}\,\negthickspace\right]\mathclose{}=\alpha e_{1}+\beta_{1}e_{k+1}+\beta_{2}e_{l+1}.$
Step 3: Combining Negative and Zero Components
Concatenate $x_{n}$ and $x_{z}$:
$x_{nz}=\beta_{1}e_{1}+\beta_{2}e_{l-k+1}.$
Combine $D_{n}$ and $D_{z}$ as:
$D_{nz}=\mathopen{}\left[{\begin{array}[]{cc}D_{n}&\\ &D_{z}\end{array}}\,\negthickspace\right]\mathclose{}.$
Applying Lemma B.2 to $\mathopen{}\left({D_{nz},x_{nz}}\,\negthickspace\right)\mathclose{}$, we find $A_{nz}$ such that:
$A_{nz}^{T}D_{nz}A_{nz}=D_{nz},\quad A_{nz}x_{nz}=\beta e_{1}.$
Define $A_{2}$ as:
$A_{2}=\mathopen{}\left[{\begin{array}[]{cc}I&\\ &A_{nz}\end{array}}\,\negthickspace\right]\mathclose{}.$
Since $A_{2}^{T}DA_{2}=D$, we have $A_{2}\in G$. Now:
$A_{2}A_{1}x=\alpha e_{1}+\beta e_{k+1}.$
Step 4: Aligning with Canonical Basis
Applying Lemma B.3 to $\mathopen{}\left({D,A_{2}A_{1}x}\,\negthickspace\right)\mathclose{}$ (for the case of $x^{T}Dx>0$), we find $A_{3}$ such that:
$A_{3}^{T}DA_{3}=D,\quad A_{3}\mathopen{}\left({A_{2}A_{1}x}\,\negthickspace\right)\mathclose{}=\gamma e_{1}.$ (44)
Similar steps hold in the case of $x^{T}Dx<0$, i.e., we can find $A_{3}$ such that $A_{3}^{T}DA_{3}=D\;\;\And\;A_{3}\mathopen{}\left({A_{2}A_{1}x}\,\negthickspace\right)\mathclose{}=\gamma e_{k+1}$. Since $A_{1},A_{2},A_{3}\in G$, their product $A_{3}A_{2}A_{1}\in G$. Let, $W=A_{3}A_{2}A_{1}$. Then, we can write,
$W^{T}DW=D,\quad Wx=\gamma e_{1}.$ (45)
Conclusion
Since, the Eq. (45) holds for any $x\in\mathbb{R}^{n}$ having the same value for $\|x\|_{D}$, we can conclude that $x,y\in\mathbb{R}^{n}$ lie in the same orbit if $\mathopen{}\left({\|x\|_{D}}\,\negthickspace\right)\mathclose{}=\mathopen{}\left({\|y\|_{D}}\,\negthickspace\right)\mathclose{}$.
See 4.2
Proof.
Since $f(x)$ is equivariant under the action of $G$, we have
$f\mathopen{}\left({g\cdot x}\,\negthickspace\right)\mathclose{}=g\cdot f(x),\quad\forall g\in G,x\in\mathbb{R}^{n}.$ (46)
This can be rewritten in terms of a canonical form as,
$f(y)=\phi_{1}(y)\cdot\phi_{2}\mathopen{}\left({r(y)}\,\negthickspace\right)\mathclose{},$ (47)
where,
$y=\phi_{1}(y)\cdot r(y).$ (48)
Here, $r(y)\in\mathbb{R}^{n}\setminus U$ is a representative element chosen from the entire orbit of $y$, denoted by $G\cdot y$. The function $\phi_{1}:\mathbb{R}^{n}\setminus U\rightarrow G$ maps each point $y$ to an element of $G$ that aligns $y$ with its orbit representative $r(y)$.
Since, $r(y)$ is the same for every element in the orbit of $y$, it follows that $\phi_{2}\circ r$ is $G$-invariant. Therefore, applying Theorem 4.1, we can express this as
$\left(\phi_{2}\circ r\right)(x)=\phi_{n}\left(\|x\|_{A}\right),$ (49)
for some function $\phi_{n}$.
Next, we select a representative element $r(y)$ such that it satisfies the homogeneity condition $r(\alpha y)=\alpha\,r(y)$ for any scalar $\alpha\in\mathbb{R}$ with $\alpha>0$. This can be achieved by setting $r(y)=\alpha e_{i}$, where $\alpha=\|y\|_{A}$ represents the norm of $y$ with respect to the matrix $A$, and the choice of the standard basis vector $e_{i}$ depends on the sign of $\|y\|_{A}$, denoted by $\operatorname{sign}\left(\|y\|_{A}\right)$.
This specific selection of $r(y)$ implies that if we define,
$\phi_{1}(\alpha y)=\phi_{1}(y),$ (50)
then the following two equations, derived from Eq. (48) and the chosen form of $r(y)$, will hold consistently:
$\alpha y=\alpha\mathopen{}\left({\phi_{1}(y)\cdot r(y)}\,\negthickspace\right)\mathclose{}$ $=\phi_{1}(y)\cdot\alpha r(y),$ (51)
$\alpha y$ $=\phi_{1}\mathopen{}\left({(\alpha y)}\,\negthickspace\right)\mathclose{}\cdot\alpha r(y).$ (52)
Thus, the Eq. (50) shows that $\phi_{1}$ is scale-invariant. Hence, we can write,
$\phi_{1}(x)=\phi_{s}\left(\frac{x}{\|x\|_{A}}\right),$ (53)
for some function $\phi_{s}$.
Finally, substituting Eq. (49) and (53) into Eq. (47), we obtain,
$f(x)=\phi_{s}\mathopen{}\left({\frac{x}{\|x\|_{A}}}\,\negthickspace\right)\mathclose{}\cdot\phi_{n}\mathopen{}\left({\|x\|_{A}}\,\negthickspace\right)\mathclose{}.$
See 4.1
Proof.
From Theorem 4.2, we know that $\phi_{s}(.)$ belongs to the group $G$. Thus, we can take the inverse (modulo the stabilizer subgroup of $x$) of $\phi_{s}$ and express the equation as follows:
$\mathopen{}\left({\phi_{s}\mathopen{}\left({\frac{x}{\|x\|_{A}}}\,\negthickspace\right)\mathclose{}}\,\negthickspace\right)\mathclose{}^{-1}f(x)=\phi_{n}\mathopen{}\left({\|x\|_{A}}\,\negthickspace\right)\mathclose{}$ (54)
If $A$ is either a positive definite or a negative definite matrix, the map $x\mapsto\frac{x}{\|x\|_{A}}$ is smooth on the domain $\mathbb{R}^{n}\setminus U$. Since the group action is smooth, we conclude that the map $x\mapsto\phi_{s}\mathopen{}\left({\frac{x}{\|x\|_{A}}}\,\negthickspace\right)\mathclose{}^{-1}$ is smooth. Therefore, if $f$ is smooth, $\phi_{n}$ must also be smooth.
In the case where $A$ is neither positive definite nor negative definite, the map $x\mapsto\frac{x}{\|x\|_{A}}$ is not smooth, but it is continuous. Hence, if $f$ is continuous, $\phi_{n}$ is continuous as well. ∎
Theorem B.4 (Classification of $p$-tuples up to the $A$-orthogonal group, cf. [Jacobson, 1985, Chapter 6]).
] Let $A$ be a non-degenerate real symmetric $n\times n$ matrix, and consider the bilinear form $\langle x,y\rangle_{A}=x^{\top}A\,y$ on $\mathbb{R}^{n}$. Suppose $(x_{1},\dots,x_{p})$ and $(y_{1},\dots,y_{p})$ are two ordered $p$-tuples in $\mathbb{R}^{n}$. Then the following are equivalent:
  1. 1.
    They have the same $A$-Gram matrix, i.e.
    $x_{i}^{\top}A\,x_{j}\;=\;y_{i}^{\top}A\,y_{j}\quad\text{for all }1\leq i,j\leq p.$
  2. 2.
    There exists an invertible linear map $Q\in\mathrm{GL}(n,\mathbb{R})$ such that
    $Q^{\top}A\,Q\;=\;A\quad\text{and}\quad Q\,x_{i}\;=\;y_{i}\quad\text{for all }i.$
Equivalently, $(x_{1},\dots,x_{p})$ and $(y_{1},\dots,y_{p})$ lie in the same orbit under the action of the group
$G\;=\;\{\,Q\in\mathrm{GL}(n,\mathbb{R}):Q^{\top}A\,Q=A\}.$
See 4.3
Proof.
In the proof of Theorem 4.2, we utilized the orbit correspondence result (Theorem 4.1) to establish the norm invariance part ($\phi_{n}$) of the function decomposition. In the setting of the diagonal action on tuples, the invariance part ($\phi_{n}$) of the decomposition follows immediately from Theorem B.4. The remainder of the proof proceeds analogously to that of Theorem 4.2. ∎